OpenStudy (zubhanwc3):

what is the area under y=x+sinx [0,pi]

OpenStudy (zubhanwc3):

\[y=x+\sin x\]

OpenStudy (zubhanwc3):

\[\left[ 0,\pi \right]\]

OpenStudy (tkhunny):

Have you considered an integral? Everything is positive on \([0,\pi]\).

OpenStudy (perl):